Ohmic status stays on active Warp v284

Good day.

Small issue here…since I have updated my warp to version 284 for the laser and did not use plasma cutter.

So today I needed the plasma cutter but the Ohmic stays on Active.

I had this only happened since the machine was new and to me there was a metal stuck in the torch head.

But now I cleaned it, removed the wire and also disconnected wire to the black box and it stays on making it unusable…is there anything I should look at?

Thank you!

1 Like

Quick question: Mach-3 or Mach-4?

I am running Mach 4.

Have you tried 283? It was a more tested release, and I don’t think there is any new stuff for the laser in 284 vs. 283. I’m not sure 284 was really meant to be released either. Its fine for laser work, but possibly not fully tested in other areas like plasma (which I don’t have).

The warp9 forum is the best place to ask about that, they are very good at responding to questions there. Index

Reverted to v283 still no go.
If I start a log in mach4 and remove the wire going to the black box shutting it down
it does register ID161 High to low. connect back ID161 Low to high.

I did email avid hopefully this can be resolved or try the floating head to see what it does.
I can in Mach4 start shut down the plasma, it read the torch voltage and everything.
cut manually with keyboard only its stuck at Ohmic active. :confused:

And the previous version of ESS plugin that you were using still works correctly?

Yes it was running fine I needed to do the update to run the laser.
It might be the laser is using a signal for the plasma :confused:

I am going to revert it all however is there a way to compare all the current settings to the change ones to atleast know where the issue can be? In the end I need to run everything…spindle, plasma , laser and rotary…I know everthing works just this I need to solve.


I know a few others who are using the laser and plasma on the same machine. I think all the plasma pins from the ESS are allocated in all of the configurations, so there shouldn’t be any conflicts if you picked an open pin for your laser gate and PWM signals (like 3-2 and 3-3). The ohmic probe is assigned to port 2 on my machine, so there shouldn’t be a conflict.

When I want to compare setup/configurations, I get the Machine.ini file from each one and load them into Notepad++ and run a compare to see whats different. Pretty much every config setting in Mach4 is in that file.


Did you end up getting to the bottom of the issue here?

After a few tests with the instruction from Avid it seems that the Ohmic black box is bad.
waiting for a replacement.

Did you go back and restore the backup of the full installation folders and see if it was failing there? It seems really ‘odd’ that that unit would just stop working out of the blue?

I have reinstalled due to Mach 4 Ghosts but nothing.
I got and replaced the board and it fixed the issue.